The sum of two number is 30 three times the first number added to twice the second number is 64. Find the number
Answer by oberobic(2304)

x + y = 30
so
x = 30-y
and
y = 30-x
.
3*x + 2*y = 64
.
substitute y = 30-x
.
3x + 2(30-x) = 64
.
3x +60 -2x = 64
.
x = 4
.
y = 30 - x
y = 30 - 4
y = 26
.
We know the sum = 30 since we used that to find 'y'.
But is the other relationship true, too?
.
3*x + 2*y = 64
3(4) + 2(26) = 12 + 52 = 64
Correct.
.
Answer: The two numbers are 4 and 26.
.
Done.